OpenStudy (anonymous):

Select the exclusion that fits best with this problem. 13x^6/51x^4 x = 0 x, y, z = 0 1 + y + y2, y = 0, 1 - y = 0 a, b = 0, a2+ ab + b2= 0 2x - 1 = 0, 4x2+ 2x + 1 = 0

OpenStudy (e.mccormick):

OK, for \[\frac{13x^6}{51x^4}\]What could make it into an invalid answer?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

and hi again :D

OpenStudy (e.mccormick):

Hi again. "Select the exclusion" means things that would not give valid answers.

OpenStudy (anonymous):

ths one: a, b = 0, a2+ ab + b2= 0?

OpenStudy (e.mccormick):

Well, there is a huge clue in the fact that it does not have an a or b in it... You have a fraction there, or rational equation. What breaks any fraction? The one type of division that they say is a big no-no, can't be done, does not exist... and so on.

OpenStudy (anonymous):

as always, you confuse me, lol

OpenStudy (e.mccormick):

LOL. Well, how about this: \[\frac{2}{0}\]Does that mean anyhting in math?

OpenStudy (e.mccormick):

Just try it in a calculator. See what \(2\div 0\) gets you.

OpenStudy (anonymous):

you can't divide by zero....it will open a wormhole to the gamma quadrant or something

OpenStudy (anonymous):

the exclusion rules: no dividing by zero no negative roots....\[\sqrt{x}\] for x>=0 ln x? x has to be greater than zero

OpenStudy (e.mccormick):

But kantalope, wormholes are fun! They het you warp space time and all that!

OpenStudy (anonymous):

o-o back and again, whaaaaat?

OpenStudy (anonymous):

oh, yeah complex infinity when u divide 2 by 0

OpenStudy (e.mccormick):

Yah. So, when the bottom would be 0 it is mathematically bad. That is what they want to know.

OpenStudy (e.mccormick):

When \(\frac{13x^6}{51x^4}=\frac{13x^6}{0}\) it is to be excluded. So, which answer does that?

OpenStudy (anonymous):

x=0?

OpenStudy (e.mccormick):

Yep!

OpenStudy (anonymous):

helpful as always, lol

OpenStudy (e.mccormick):

Well... I poke, I prod... I try not to do more than clues. Sometimes it works, sometimes not.

OpenStudy (anonymous):

lol, i think your tactic is confuse then answer

OpenStudy (e.mccormick):

No, no. Math is progressive. So I dig up old topics like 2/0 to indicate how this relates to the other topic!

OpenStudy (anonymous):

oh so it's confuse, help confuse now? xD

OpenStudy (e.mccormick):

It is called the Socratic Method. A tried and true form of teaching!

OpenStudy (anonymous):

...so I am right :D

OpenStudy (e.mccormick):

Well, because there is an end goal, it is not pure Socratic Method, but the baseline idea is there. Use questions to expose the core of the problem and related it to previously learned facts.

OpenStudy (anonymous):

......yes